# TokenTrace: Multi-Concept Attribution through Watermarked Token Recovery

### v1

@@ 2 Related Works
Proactive schemes embed signals into the generative process for tasks like deepfake detection [42, 52], manipulated content localization [3, 49], and concept attribution [1, 2]. While pixel-domain methods are fragile and fail when multiple concept spatially overlap, and latent-domain methods are robust but content-agnostic, a new category operates in the semantic domain [26], linking watermarks to textual representations. For instance, ”Guarding Textual Inversion” [12] is designed for single-concept tracing and does not address compositional attribution. CustomMark [2] also customizes models via text prompts for multi-concept attribution but struggles to disentangle concepts that are spatially overlapping. Thus, a robust, query-based mechanism for disentangling composed concepts remains a significant and unaddressed challenge.
@@ 3 Method
In this section, we present our method, TokenTrace, a proactive watermarking framework designed for robust, multi-concept attribution. As illustrated in Figure 2, our framework is composed of two primary stages: (a) a concept encoding stage that embeds concept secrets into the generative process, (b) a concept decoding stage that retrieves the corresponding concept secrets that indicated by a query prompt, and (c) the detailed architecture of our TokenTrace module.
@@ 3.1 Concept Encoding
  • The concept encoder ($f_{enc}$) targets the specific concept token to be watermarked. A user prompt embedding $E_{prompt}$ is a collection of token embeddings $\{e_{1},\dots,e_{c},\dots,e_{k}\}$, where $e_{c}$ is the target concept token embedding. The encoder $f_{enc}$ takes the concept secret ($\mathcal{S}$) and the corresponding concept token embedding ($e_{c}$) as input to generate a perturbation. This perturbation is fused (via element-wise addition) only with the target token embedding $e_{c}$ to create a perturbed token embedding $\hat{e}_{c}$. The final perturbed prompt embedding, $\hat{E}_{prompt}$, is calculated by replacing the original token:
    $\left\{\begin{aligned} &\hat{e}_{c}=e_{c}+f_{enc}(e_{c},\mathcal{S})\\ &\hat{E}_{prompt}=\{e_{1},\dots,\hat{e}_{c},\dots,e_{k}\}\end{aligned}\right$. (1)
  • The secret mapper ($f_{map}$) takes only the concept secret ($\mathcal{S}$) as input and generates a structured perturb gaussian pattern. This pattern is fused with the original noise pattern ($z_{T}$) to create the perturbed initial noise, $\hat{z}_{T}$.
    $\hat{z}_{T}=z_{T}+f_{map}(\mathcal{S})$ (2)
Finally, the diffusion model ($DM$) is conditioned on both of these inputs, the prompt embedding with perturbed concept token embeddings ($\hat{E}_{prompt}$) and the perturbed noise ($\hat{z}_{T}$), to generate the final watermarked image ($I_{wm}$).
$I_{wm}=DM(\hat{z}_{T},\hat{E}_{prompt})$ (3)
@@ 3.2 Concept Decoding
As shown in Figure 2(b), the decoding stage is a two-step pipeline designed to reliably retrieve the embedded watermark and confirm its origin. First, the watermarked image ($I_{wm}$) and a textual query ($P_{query}$) are fed into the TokenTrace module ($f_{tt}$). This textual query is a simple prompt, selected from a pre-defined set, that specifies which concept to retrieve. For example, to retrieve the <<sks-object>> concept, the query $P_{query}$ would be ”a photo of <<sks-object>>”. The TokenTrace module then uses both inputs to predict the concept embedding ($\tilde{e}_{c}$).
$\tilde{e}_{c}=f_{tt}(I_{wm},P_{query})$ (4)
This predicted embedding is then passed to the secret decoder ($f_{dec}$), which we implement as a simple linear network. This decoder translates the high-dimensional embedding back into the original bit-secret, $\tilde{\mathcal{S}}$.
$\tilde{\mathcal{S}}=f_{dec}(\tilde{e}_{c})$ (5)
This query-based pipeline provides a direct solution to the multi-concept challenge. By requiring a specific text query ($P_{query}$) to initiate retrieval, the system can selectively isolate and verify a single concept’s signature. This mechanism allows it to successfully disentangle and attribute multiple, overlapping concepts (e.g., an object and a style) from a single image, a task that pixel-based methods cannot perform.
@@
    The watermarked image $I_{wm}$ is passed through the frozen Image encoder ($f_{img\_enc}$). Its output features are then fed into a trainable projection layer ($f_{proj_{1}}$) to align their dimensionality with the text features.
@@
    The query prompt $P_{query}$ is passed through the frozen Text encoder ($f_{text\_enc}$).
@@
    The aligned image features and the text features are fed into a trainable attention module ($f_{attn}$), which produces a fused, context-aware representation.
@@
    This fused representation is passed through a final trainable projection layer ($f_{proj_{2}}$) to generate the predicted concept embedding, $\tilde{e}_{c}$.
@@
$\left\{\begin{aligned} &F_{img}=f_{proj_{1}}(f_{img\_enc}(I_{wm}))\\ &F_{text}=f_{text\_enc}(P_{query})\\ &F_{fused}=f_{attn}(F_{img},F_{text})\\ &\tilde{e}_{c}=f_{proj_{2}}(F_{fused})\end{aligned}\right$. (6)
@@ 3.3 Training Objective
Our training objective is to jointly optimize all trainable components: the concept encoder, secret mapper, the trainable projection and attention layers of the TokenTrace module, and the secret decoder. The model is trained using a composite loss function, $\mathcal{L}_{total}$, which is designed to simultaneously satisfy two competing goals: watermark retrieval accuracy and visual fidelity. This final loss is a weighted sum of four individual constraints:
@@
    $\mathcal{L}_{CE}=\text{BCE}(\mathcal{S},\sigma(\hat{\mathcal{S}}_{logits})),$ (7)
    where $\sigma$ is the sigmoid function, $\mathcal{S}$ is the ground-truth, and $\hat{\mathcal{S}}_{logits}$ is the output logits from the secret decoder.
@@
    $\mathcal{L}_{CSD}=1-\frac{\phi(I_{clean})\cdot\phi(I_{wm})}{\|\phi(I_{clean})\|\|\phi(I_{wm})\|},$ (8)
    where $\phi$ represents the feature extractor, $I_{clean}$ is the clean image, and $I_{wm}$ is the watermarked image.
@@
    $\mathcal{L}_{L2}=\|I_{clean}-I_{wm}\|_{2}^{2}.$ (9)
@@
    $\mathcal{L}_{reg}=\|e_{c}-\tilde{e}_{c}\|_{2}^{2},$ (10)
    where $e_{c}$ and $\tilde{e}_{c}$ are the ground truth and predicted concept embeddings.
@@
$\mathcal{L}_{total}=\lambda_{1}\mathcal{L}_{CE}+\lambda_{2}\mathcal{L}_{CSD}+\lambda_{3}\mathcal{L}_{L2}+\lambda_{4}\mathcal{L}_{reg},$ (11)
where $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ are the trade-off parameters to balance different kind of constrains. By using this composite loss function, we can jointly optimize for two competing objectives: retrieval accuracy (via $\mathcal{L}_{CE}$ and $\mathcal{L}_{reg}$) and visual fidelity (via $\mathcal{L}_{L2}$ and $\mathcal{L}_{CSD}$). This ensures the resulting watermark is both robustly detectable and imperceptible.
The end-to-end training procedure for TokenTrace is detailed in Algorithm 1 in the supplementary material. In summary, we jointly optimize all trainable parameters. In each training iteration, we sample a batch of data (including a clean image, prompts, secret, and ground-truth embedding) and perform the concept encoding step to generate a watermarked image. This image is then passed through the decoding pipeline to retrieve the secret. Finally, we compute the total composite loss $\mathcal{L}_{total}$ and update all trainable parameters via gradient descent. The inference (verification) process is a simple forward pass of the trained decoding modules ($f_{tt}$ and $f_{dec}$), as described in Section 3.2.
@@ 4.1 Datasets
For the single style concept attribution task, we employ $23$ distinct styles sourced from the WikiArt dataset [37], which is a collection of digital artworks sourced from the online art encyclopedia  [^2] https://www.wikiart.org/. It is one of the most popular datasets for art-related computer vision tasks. For the single object concept attribution task, we use $1000$ object classes from the ImageNet dataset [10], which is a foundational, large-scale dataset in computer vision. For our primary task of customized multi-concept attribution, we source pre-trained concept embeddings from the Stable Diffusion Textual Inversion Embeddings library  [^3] https://cyberes.github.io/stable-diffusion-textual-inversion-models/. This allows us to construct challenging prompts that combine multiple concepts, such as a specific object and a specific style, to test our model’s disentanglement capabilities. We also evaluate general multi-concept attribution, as composing multiple customized textual inversion embeddings in a single prompt often degrades image quality. To isolate our attribution mechanism’s performance from this generation artifact, we use standard, non-customized concepts (e.g., ”a cat and a dog”). We construct a new evaluation benchmark by using ChatGPT [6] to generate a diverse set of prompts that involve the composition of multiple concepts. For each prompt, we also generate a list of target concepts that are presented in the prompt  [^4] Specific examples can be found in supplementary material..
@@ 4.2 Experimental Settings
Table 1: Comparison of attribution performance on the WikiArt and ImageNet datasets. TokenTrace achieves the highest accuracy across both datasets, outperforming all passive and proactive baselines. ”Bit” represents bit accuracy (%) and ”Att.” represents attribution accuracy (%). Passive baselines do not have Bit Accuracy as they do not retrieve a secret bit-string.
@@
Figure 3: Qualitative analysis of visual fidelity for watermarked images. (a) Results on abstract artistic style concepts from the WikiArt dataset. (b) Results on diverse object concepts from the ImageNet dataset. In both cases, the ”Watermarked” images are visually indistinguishable from the ”Clean” originals.
Table 2: Performance comparison for multi-concept attribution. The left plot shows results for attributing 2 customized concepts (one object and one style), while the right plot shows results for 4 general concepts. TokenTrace, consistently outperforms the CustomMark baseline in both bit accuracy, represented as ”Bit”, and attribution accuracy, represented as ”Attrbution”. Performance is further improved by applying prompt weighting (TokenTraceP). All accuracy metrics are in percent (%).
@@ 4.3 Results and Analysis
The qualitative results in Figure 3(a) further validate our method’s performance. The ”Watermarked” images successfully retain the intricate details and overall aesthetic of the original artistic styles. The visual fidelity between the ”Clean” and ”Watermarked” pairs is high. This confirms that watermarks inserted by TokenTrace is imperceptible, a crucial requirement for any practical attribution system.
@@
We further assess evaluate the scalability and performance of TokenTrace on discrete objects, we conducted a large-scale experiment on $1000$ object concepts sourced from the ImageNet [10] dataset. Similar from the style attribution task, the results presented in Table 1 clearly demonstrates the superiority of our approach. TokenTrace achieves a state-of-the-art attribution accuracy of $90.43\%$ and a bit accuracy of $95.82\%$. It significantly outperforms all passive methods and shows a clear performance gain over the other proactive baselines, ProMark ($87.30\%$) and CustomMark ($87.12\%$). The qualitative results in Figure 3(b) further confirm the imperceptibility of our watermark. Across diverse concepts, the ”Watermarked” images are visually indistinguishable from their ”Clean” images, demonstrating that TokenTrace achieves robust attribution while maintaining high visual fidelity.
@@
Finally, Figure 4 and Figure 5 provides qualitative proof of this disentanglement. Specifically, from a single image generated with four watermarked concepts (e.g., ”detailed cat wearing a sweater… glowing rays”), our query-based module successfully retrieves the correct, independent secret for each concept, retrieving ’cat’ and ’sweater’ with $100\%$ accuracy and ’detailed’ and ’rays’ with high accuracy. This confirms our query-based mechanism solves concept overlap and scales to complex, compositional prompts. More qualitative examples can be found in the supplementary material.
@@ 4.4 Ablation Studies
Our composite loss function ($\mathcal{L}_{total}$), as defined in Eq. 11, is designed to jointly optimize for retrieval accuracy and visual fidelity. To demonstrate the contribution of each loss term, we train our model from scratch with different components of the objective function removed. The full model (”All”) is trained with all loss components, including $\mathcal{L}_{CE}$, $\mathcal{L}_{CSD}$, $\mathcal{L}_{L2}$, and $\mathcal{L}_{reg}$. As shown in Table 3, our complete model (”All”) achieves the best performance. Removing the CSD loss (”No CSD”) causes the most significant performance drop. This finding confirms that the $\mathcal{L}_{CSD}$ term is critical for maintaining the high-level semantic and style consistency required for accurate retrieval. Removing the image-space L2 loss (”No L2 (Image)”) also leads to a notable decrease in performance, with attribution accuracy dropping to $86.37\%$ and the CSD Score falling to $0.73$. The latent-space L2 (”No L2 (Latent)”) also contributes, with its removal dropping accuracy to $88.52\%$.
Table 3: Ablation study for the objective function on the WikiArt dataset. Our full model (”All”) significantly outperforms variants where individual loss components (CSD, Latent L2, Image L2) are removed, demonstrating that each component is necessary for achieving both high attribution accuracy and visual fidelity.
@@
Table 4: Ablation study for different types of distortions on the WikiArt dataset. Our method demonstrates high robustness, maintaining strong attribution accuracy even after significant image transformations.

### v2

@@ 2 Related Works
Proactive schemes embed signals into the generative process for tasks like deepfake detection [40, 50], manipulated content localization [3, 47, 4], and concept attribution [1, 2]. While pixel-domain methods are fragile and fail when multiple concept spatially overlap, and latent-domain methods are robust but content-agnostic, a new category operates in the semantic domain [24], linking watermarks to textual representations. For instance, “Guarding Textual Inversion” [12] is designed for single-concept tracing and does not address compositional attribution. CustomMark [2] also customizes models via text prompts for multi-concept attribution but struggles to disentangle concepts that are spatially overlapping. Thus, a robust, query-based mechanism for disentangling composed concepts remains a significant and unaddressed challenge.
@@ 3 Method
In this section, we present our method, TokenTrace, a proactive watermarking framework designed for robust, multi-concept attribution that can be used as a controlled-generation provenance tool. As illustrated in Figure 2, our framework is composed of two stages: (a) a concept encoding stage that embeds concept secrets into the generative process, (b) a concept decoding stage that retrieves the concept secrets that is indicated by a query prompt, and (c) the detailed architecture of our TokenTrace module.
@@ 3.1 Concept Encoding
  • The concept encoder ($f_{\text{enc}}$) targets the specific concept token to be watermarked. A user prompt embedding $E_{\text{prompt}}$ is a collection of token embeddings $\{e_{\text{1}},\dots,e_{\text{c}},\dots,e_{\text{k}}\}$, where $e_{\text{c}}$ is the target concept token embedding. The encoder $f_{\text{enc}}$ takes the concept secret ($\mathcal{S}$) and the corresponding concept token embedding ($e_{\text{c}}$) as input to generate a perturbation. This perturbation is fused (via element-wise addition) only with the target token embedding $e_{\text{c}}$ to create a perturbed token embedding $\hat{e}_{\text{c}}$. The final perturbed prompt embedding, $\hat{E}_{\text{prompt}}$, is calculated by replacing the original token:
    $\left\{\begin{aligned} &\hat{e}_{\text{c}}=e_{\text{c}}+f_{\text{enc}}(e_{\text{c}},\mathcal{S})\\ &\hat{E}_{\text{prompt}}=\{e_{\text{1}},\dots,\hat{e}_{\text{c}},\dots,e_{\text{k}}\}\end{aligned}\right$. (1)
  • The secret mapper ($f_{\text{map}}$) takes only the concept secret ($\mathcal{S}$) as input and generates a structured perturb gaussian pattern. This pattern is fused with the original noise pattern ($z_{\text{T}}$) to create the perturbed initial noise, $\hat{z}_{\text{T}}$.
    $\hat{z}_{\text{T}}=z_{\text{T}}+f_{\text{map}}(\mathcal{S})$ (2)
Finally, the diffusion model ($DM$) is conditioned on both of these inputs, the prompt embedding with perturbed concept token embeddings ($\hat{E}_{\text{prompt}}$) and the perturbed noise ($\hat{z}_{\text{T}}$), to generate the final watermarked image ($I_{\text{wm}}$).
$I_{\text{wm}}=DM(\hat{z}_{\text{T}},\hat{E}_{\text{prompt}})$ (3)
@@ 3.2 Concept Decoding
As shown in Figure 2(b), the decoding stage is a two-step pipeline designed to reliably retrieve the embedded watermark and confirm its origin. First, the watermarked image ($I_{\text{wm}}$) and a textual query ($P_{\text{query}}$) are fed into the TokenTrace module ($f_{\text{tt}}$). This textual query is a simple prompt, selected from a pre-defined set, that specifies which concept to retrieve. For example, to retrieve the <<sks-object>> concept, the query $P_{\text{query}}$ would be “a photo of <<sks-object>>”. The TokenTrace module then uses both inputs to predict the concept embedding ($\tilde{e}_{\text{c}}$).
$\tilde{e}_{\text{c}}=f_{\text{tt}}(I_{\text{wm}},P_{\text{query}})$ (4)
This predicted embedding is then passed to the secret decoder ($f_{\text{dec}}$), which we implement as a simple linear network. This decoder translates the high-dimensional embedding back into the original bit-secret, $\tilde{\mathcal{S}}$.
$\tilde{\mathcal{S}}=f_{\text{dec}}(\tilde{e}_{\text{c}})$ (5)
This query-based pipeline provides a direct solution to the multi-concept challenge. By requiring a specific text query ($P_{\text{query}}$) to initiate retrieval, the system can selectively isolate and verify a single concept’s signature. This mechanism allows it to successfully disentangle and attribute multiple, overlapping concepts (e.g., an object and a style) from a single image, a task that pixel-based methods cannot perform.
@@
    The watermarked image $I_{\text{wm}}$ is passed through the frozen Image encoder ($f_{\text{img\_enc}}$). Its output features are then fed into a trainable projection layer ($f_{\text{proj\_1}}$) to align their dimensionality with the text features.
@@
    The query prompt $P_{\text{query}}$ is passed through the frozen Text encoder ($f_{\text{text\_enc}}$).
@@
    The aligned image features and the text features are fed into a trainable attention module ($f_{\text{attn}}$), which produces a fused, context-aware representation.
@@
    This fused representation is passed through a final trainable projection layer ($f_{\text{proj\_2}}$) to generate the predicted concept embedding, $\tilde{e}_{\text{c}}$.
@@
$\left\{\begin{aligned} &F_{\text{img}}=f_{\text{proj\_1}}(f_{\text{img\_enc}}(I_{\text{wm}}))\\ &F_{\text{text}}=f_{\text{text\_enc}}(P_{\text{query}})\\ &F_{\text{fused}}=f_{\text{attn}}(F_{\text{img}},F_{\text{text}})\\ &\tilde{e}_{\text{c}}=f_{\text{proj\_2}}(F_{\text{fused}})\end{aligned}\right$. (6)
@@ 3.3 Training Objective
Our training objective is to jointly optimize all trainable components: the concept encoder, secret mapper, the trainable projection and attention layers of the TokenTrace module, and the secret decoder. The model is trained using a composite loss function, $\mathcal{L}_{\text{total}}$, which is designed to simultaneously satisfy two competing goals: watermark retrieval accuracy and visual fidelity. This final loss is a weighted sum of four individual constraints:
@@
    $\mathcal{L}_{\text{BCE}}=\text{BCE}(\mathcal{S},\sigma(\hat{\mathcal{S}}_{\text{logits}})),$ (7)
    where $\sigma$ is the sigmoid function, $\mathcal{S}$ is the ground-truth, and $\hat{\mathcal{S}}_{\text{logits}}$ is the output logits from the secret decoder.
@@
    $\mathcal{L}_{\text{CSD}}=1-\frac{\phi(I_{\text{clean}})\cdot\phi(I_{\text{wm}})}{\|\phi(I_{\text{clean}})\|\|\phi(I_{\text{wm}})\|},$ (8)
    where $\phi$ represents the feature extractor, $I_{\text{clean}}$ is the clean image, and $I_{\text{wm}}$ is the watermarked image.
@@
    $\mathcal{L}_{\text{L2}}=\|I_{\text{clean}}-I_{\text{wm}}\|_{2}^{2}.$ (9)
@@
    $\mathcal{L}_{\text{reg}}=\|e_{\text{c}}-\tilde{e}_{\text{c}}\|_{2}^{2},$ (10)
    where $e_{\text{c}}$ and $\tilde{e}_{\text{c}}$ are the ground truth and predicted concept embeddings.
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{BCE}}+\lambda_{2}\mathcal{L}_{\text{CSD}}+\lambda_{3}\mathcal{L}_{\text{L2}}+\lambda_{4}\mathcal{L}_{\text{reg}},$ (11)
where $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ are the trade-off parameters to balance different kind of constrains. By using this composite loss function, we can jointly optimize for two competing objectives: retrieval accuracy (via $\mathcal{L}_{\text{BCE}}$ and $\mathcal{L}_{\text{reg}}$) and visual fidelity (via $\mathcal{L}_{\text{L2}}$ and $\mathcal{L}_{\text{CSD}}$). This ensures the resulting watermark is both robustly detectable and imperceptible.
The end-to-end training procedure for TokenTrace is detailed in Algorithm 1 in the supplementary material. In summary, we jointly optimize all trainable parameters. In each training iteration, we sample a batch of data (including a clean image, prompts, secret, and ground-truth embedding) and perform the concept encoding step to generate a watermarked image. This image is then passed through the decoding pipeline to retrieve the secret. Finally, we compute the total composite loss $\mathcal{L}_{\text{total}}$ and update all trainable parameters via gradient descent. The inference (verification) process is a simple forward pass of the trained decoding modules ($f_{\text{tt}}$ and $f_{\text{dec}}$), as described in Section 3.2.
@@ 4.1 Datasets
For the single style concept attribution task, we employ $23$ distinct styles sourced from the WikiArt dataset [35], which is a collection of digital artworks sourced from the online art encyclopedia  [^2] https://www.wikiart.org/. It is one of the most popular datasets for art-related computer vision tasks. For the single object concept attribution task, we use $1000$ object classes from the ImageNet dataset [10], which is a foundational, large-scale dataset in computer vision. For our primary task of customized multi-concept attribution, we source pre-trained concept embeddings from the Stable Diffusion Textual Inversion Embeddings library  [^3] https://cyberes.github.io/stable-diffusion-textual-inversion-models/. This allows us to construct challenging prompts that combine multiple concepts, such as a specific object and a specific style, to test our model’s disentanglement capabilities. We also evaluate general multi-concept attribution, as composing multiple customized textual inversion embeddings in a single prompt often degrades image quality. To isolate our attribution mechanism’s performance from this generation artifact, we use standard, non-customized concepts (e.g., “a cat and a dog”). We construct a new evaluation benchmark by using ChatGPT [7] to generate a diverse set of prompts that involve the composition of multiple concepts. For each prompt, we also generate a list of target concepts that are presented in the prompt  [^4] Specific examples can be found in supplementary material..
@@ 4.2 Experimental Settings
Table 1: Comparison of attribution performance on the WikiArt and ImageNet datasets. TokenTrace achieves the highest accuracy across both datasets, outperforming all passive and proactive baselines. “Bit” represents bit accuracy (%) and “Att.” represents attribution accuracy (%). Passive baselines do not have Bit Accuracy as they do not retrieve a secret bit-string.
@@
Figure 3: Qualitative analysis of visual fidelity for watermarked images. (a) Results on abstract artistic style concepts from the WikiArt dataset. (b) Results on diverse object concepts from the ImageNet dataset. In both cases, the “Watermarked” images are visually indistinguishable from the “Clean” originals.
Table 2: Performance comparison for multi-concept attribution. The left plot shows results for attributing 2 customized concepts (one object and one style), while the right plot shows results for 4 general concepts. TokenTrace, consistently outperforms the CustomMark baseline in both bit accuracy, represented as “Bit”, and attribution accuracy, represented as “Attrbution”. Performance is further improved by applying prompt weighting (TokenTraceP). All accuracy metrics are in percent (%).
@@ 4.3 Results and Analysis
The qualitative results in Figure 3(a) further validate our method’s performance. The “Watermarked” images successfully retain the intricate details and overall aesthetic of the original artistic styles. The visual fidelity between the “Clean” and “Watermarked” pairs is high. This confirms that watermarks inserted by TokenTrace is imperceptible, a crucial requirement for any practical attribution system.
@@
We further assess evaluate the scalability and performance of TokenTrace on discrete objects, we conducted a large-scale experiment on $1000$ object concepts sourced from the ImageNet [10] dataset. Similar from the style attribution task, the results presented in Table 1 clearly demonstrates the superiority of our approach. TokenTrace achieves a state-of-the-art attribution accuracy of $90.43\%$ and a bit accuracy of $95.82\%$. It significantly outperforms all passive methods and shows a clear performance gain over the other proactive baselines, ProMark ($87.30\%$) and CustomMark ($87.12\%$). The qualitative results in Figure 3(b) further confirm the imperceptibility of our watermark. Across diverse concepts, the “Watermarked” images are visually indistinguishable from their “Clean” images, demonstrating that TokenTrace achieves robust attribution while maintaining high visual fidelity.
@@
Finally, Figure 4 and Figure 5 provides qualitative proof of this disentanglement. Specifically, from a single image generated with four watermarked concepts (e.g., “detailed cat wearing a sweater… glowing rays”), our query-based module successfully retrieves the correct, independent secret for each concept, retrieving ’cat’ and ’sweater’ with $100\%$ accuracy and ’detailed’ and ’rays’ with high accuracy. This confirms our query-based mechanism solves concept overlap and scales to complex prompts. More qualitative examples can be found in the supplementary material.
@@ 4.4 Ablation Studies
Our composite loss function ($\mathcal{L}_{\text{total}}$), as defined in Eq. 11, is designed to jointly optimize for retrieval accuracy and visual fidelity. To demonstrate the contribution of each loss term, we train our model from scratch with different components of the objective function removed. The full model (“All”) is trained with all loss components, including $\mathcal{L}_{\text{BCE}}$, $\mathcal{L}_{\text{CSD}}$, $\mathcal{L}_{\text{L2}}$, and $\mathcal{L}_{\text{reg}}$. As shown in Table 3, our complete model (“All”) achieves the best performance. Removing the CSD loss (“No CSD”) causes the most significant performance drop. This finding confirms that the $\mathcal{L}_{\text{CSD}}$ term is critical for maintaining the high-level semantic and style consistency required for accurate retrieval. Removing the image-space L2 loss (“No L2 (Image)”) also leads to a notable decrease in performance, with attribution accuracy dropping to $86.37\%$ and the CSD Score falling to $0.73$. The latent-space L2 (“No L2 (Latent)”) also contributes, with its removal dropping accuracy to $88.52\%$.
Table 3: Ablation study for the objective function on the WikiArt dataset. Our full model (“All”) significantly outperforms variants where individual loss components (CSD, Latent L2, Image L2) are removed, demonstrating that each component is necessary for achieving both high attribution accuracy and visual fidelity.
@@
Table 4: Ablation study for different types of distortions on the WikiArt dataset. Our method demonstrates high robustness, maintaining strong attribution accuracy. The detailed description of “Adversarial Attack” can be found in supplementary material.